# Modified Axelrod Model Showing Opinion Convergence And Polarization In Realistic Scale-Free Networks

### v1

@@ Abstract
Axelrod model is an opinion dynamics model such that each agent on a two-dimensional lattice has a finite number of possible opinions. The model is not realistic as all agents will eventually share the same opinion on every issue. Here we identify certain limitations inherent in the Axelrod model and address them by modifying some of the model’s rules. Specifically, by considering agents in a scale-free network and incorporating a more realistic interaction rule with continuous values representing agents’ opinions, our modified model offers a more comprehensive exploration of opinion dynamics. Computer simulation of our model shows a notable trend in opinion polarization on all features in the majority of reasonable simulation parameters. To mitigate this polarization, we introduce agents that work actively to reduce opinion differences. However, our findings indicate limited success in the approach. This research contributes to the understanding of opinion dynamics within society and highlights the nuanced complexities that arise when considering factors such as network structure and continuous opinion values. Our results prompt further exploration and open avenues for future investigations into effective methods of reducing opinion polarization.
Axelrod model, opinion dynamics, opinion polarization, scale-free network
@@ I Introduction
In socio-physics, opinion dynamics of individuals or groups is modeled via interactions among them. By investigating the emergence and convergence of individual opinions into collective opinions, the socio-physical approach aims to shed light on the underlying mechanisms. In particular, sociophysicists wanted to understand the cause and evolution of opinion formation and dissemination through social influence and networks, the conditions that lead to consensus or polarization, and the factors governing the resulting distributions of opinions [1].
@@
Numerous models of opinion dynamics have been proposed throughout the years. They include the Sznajd model [2, 3], Hegselmann-Krause model [4], Galam models [5], Axelrod model [6], Deffuant-Weisbuch model [7, 8] and their variants [9, 10, 11, 12, 13]. Our research here focuses on the Axelrod model [6]. In this model, individuals that we call agents from now on are located on a two-dimensional lattice. The opinion of each agent evolves by interacting with neighboring agents according to a fixed set of rules to be described in Sec. II. In particular, Axelrod found that after a sufficiently long time, all agents on the lattice tend to share the same traits for all features. That is, the traits of opinion (or opinion for short) in society converge [6]. Thus, this model fails to capture the fragmentation of cultures or political stances in numerous countries and societies in our contemporary world.
Various groups of social scientists have proposed reasons for this fragmentation in the real world [14, 15, 16, 17, 18]. For example, Ashraf and Galor [14] argued that genetic diversity, shaped by the migration of humans out of Africa, is a fundamental determinant of cultural heterogeneity. Dandekar et al. [15] tried to explain the polarization by opinion formation. Socio-physical models could offer a complementary approach to understanding the dynamics of opinion fragmentation. Different scholars have tried different simulation models attempting to capture the fragmentation behaviors in society. For instance, Chau et. al. proposed a model for opinion formation based on the social judgment theory extended from the Jager-Amblard model [19, 13]. That model incorporates assimilation and contrast, which influence agents’ opinion acceptance or rejection. More importantly, it shows opinion clustering around extremes and/or moderate. Furthermore, with slow variation of model parameters, Chau et al. observed first-order phase transition in the distribution of opinion [13].
@@
Could the Axelrod model be modified to give fragmented opinions or to incorporate cultural diversity in the long time limit? Various authors have addressed this issue over the years with partial success. For example, MacCarron et al. [20] considered a more complicated interaction rule and analyzed the clustering behaviors. They showed that by decreasing a certain tunable parameter of the model known as the agreement threshold, agents on the lattice will no longer be homogeneous in the sense that they have more than one set of features in the long run. Along another line, Battiston et al. [21] proposed that the interactions in real society are carried out in multiple layers, which correspond to different interests or topics. By considering this multiplicity in their model, they qualitatively presented new dynamical behavior that leads to the emergence and stability of cultural diversity final state [21]. Worth noticing that both MacCarron et al. and Battiston et al. considered discrete feature traits. Along a different line, Amorim modified the Axelrod model by aligning it with Bourdieu’s theory on the relation between an individual’s taste and the social environment [22].
This paper introduces a socio-physical model of opinion dynamics that builds upon the Axelrod model, incorporating continuous attitude change among agents within a scale-free network. Through analysis, our model demonstrates that agents within society gradually develop strong opinions based on our evolutionary rules. These findings are consistent with observed social behavior and hence give novel insights into the factors influencing collective human opinions. In addition, we attempted to solve the polarization problem that is so urgently needed in our currently highly divided and politicized society by modifying our model. Specifically, we set a small portion of agents who actively try to reduce opinion differences. Nonetheless, this approach can only reduce opinion polarization slightly. This work is an extension of a short project done by XZ under the supervision of HFC, the teacher of the course “Data Analysis And Modeling In Physics” [23].
@@ II The Axelrod Model
One of the most frequently used ways of measuring individual opinions in a society is to collect data from attitudinal surveys over a long period of time. It is instructive to model the attitude evolution of individuals and interest groups. One such model was proposed by Axelrod in 1997 [6]. In the Axelrod model, agents are placed on a $N\times N$ square lattice with closed boundary conditions. An agent is conveniently labeled by its position $(i,j)$ on the lattice. Each agent is characterized by the same fixed set of $F$ features =$\mathcal{F}=\{f_{1},f_{2},\cdots,f_{F}\}$. Each feature takes only values from a fixed set of traits. For simplicity, each $f_{i}\in\mathcal{F}$ is taken from the same set of $q$ traits labeled $\{0,1,\dots,q-1\}$. As a result, each agent has $F^{q}$ possible internal states. For example, one of the features could be the primary language spoken by the agent, and the corresponding set of traits is the set of all contemporary languages. Another possible feature is one’s religious belief. Each feature of an agent is randomly assigned in the beginning. Furthermore, the system evolves for several time steps as follows.
@@
Figure 1: Typical time evolution of 100 agents on a $10\times 10$ lattice with $F=5$ and $q=10$ according to the Axelrod model [6]. Adjacent agents are connected by different line styles according to their cultural similarities $F_{\text{comm}}/F$ (solid for $F_{\text{comm}}/F\leq 20\%$, dashed-dotted for $40\%$, dotted for $60\%$, dashed for $80\%$ and no line for $100\%$. It shows the convergence of opinions over time.
The system evolves according to the above rules repeatedly and its eventual behavior is analyzed. Fig. 1 reproduces the opinion convergence behavior of the Axelrod model using parameters in their paper, namely, for $10\times 10$ nodes on a lattice with $F=5$ and $q=10$ [6].
Worth mentioning is that throughout the years, many opinion dynamics models have been proposed. In particular, Deffuant et al. put forward several agent-based models with continuous opinion change such that the involved agents have limited ranges of acceptance and non-commitment, where only the impact of assimilation is taken into account [8, 7].
@@ III Deficiencies Of The Axelrod Model And Our Remedy
The Axelrod model is simple and useful in opinion dynamics. However, it has several deficiencies. We point them out and report ways to modify the model to make it more realistic below.
@@ III.1 Network Topology
In the Axelrod model, agents form a two-dimensional lattice. This network topology is artificial. More importantly, it is not compatible with the fact that many social structures can be described by scale-free networks [24]. In simple terms, all but those agents in the lattice boundary in the Axelrod model have equal strength in influencing others or being influenced by others. In contrast, scale-free network topology allows the existence of key opinion leaders that are so commonly found in the real world. Because of these, we propose putting agents in an (undirected) scale-free network in which only neighboring agents are allowed to interact directly. For simplicity, we generate such a network using the Barabási–Albert (B-A) model [25] as it is one of the most studied and used scale-free networks. Besides, it has the additional advantage that it can be efficiently generated. In this study, we simplify matters by considering non-evolving networks. It is instructive to consider evolving networks in future studies.
@@ III.2 Feature Traits
In the Axelrod model, every agent is described by a set of features, and each feature can have several possible discrete traits [6]. While there are features that are naturally described by discrete traits, the Axelrod model is incomplete as traits of a feature need not be discrete. For example, one of the features could be the agent’s fluency in a particular spoken language. Here, the possible traits are more naturally described by a continuous real number than discrete states. More importantly, using continuous traits is a natural platform to model the possibly increasingly polarized opinion distribution. We, therefore, modify the model by replacing the discrete set of traits with a real number representing the strength of the agent’s feature. We also call this type of continuous feature traits opinions of the agent.
How could we take care of the increasingly polarized opinion distribution? The naive way of representing opinions by a real number may not work too well owning to rounding errors and numerical instabilities in opinion evolution. Our resolution is to borrow the idea of the bounded confidence model [7, 4, 8, 19] by representing opinions using a real number in a bounded interval. Here we stress that our approach differs from the bounded confidence model in two ways. First, we restrict opinions in the open interval $(-1,1)$ instead of the closed one $[-1,1]$ so that there is always room for further opinion polarization. Second, we replace the standard Euclidean distance between two opinions $d_{\text{E}}(p,q)=|p-q|$ by the hyperbolic distance metric (of the Poincaré ball model), namely,
$d_{\text{h}}(p,q)=\ln\left(\frac{|1-pq|+|q-p|}{|1-pq|-|q-p|}\right)$ (1)
for any $p,q\in(-1,1)$. In other words, opinions are placed in a hyperbolic distance space. Note that $d_{\text{h}}$ diverges if and only if $(p,q)$ approaches $(1,-1)$ or $(-1,1)$. Hence, the hyperbolic distance metric on the interval $(-1,1)$ conveniently describes opinion polarization without significantly affecting the accuracy of our simulations. As for the initial values of each trait, instead of drawing from a uniform distribution in $(-1,1)$, we use a more realistic normal distribution with respect to the hyperbolic metric. Its detailed construction can be found in Sec. III.3 below.
@@ III.3 Interaction Rules Among Agents
Recall that agents change their traits according to rule 4 in the Axelrod model. This rule can only increase the cultural similarity between the chosen pair of agents. More importantly, numerical experiments tell us that the system becomes homogeneous eventually under this rule [6]. In other words, rule 4 is too cohesive. Besides, it is not consistent with the observation that cultural diversity among groups of people appears even in apparently culturally homogeneous societies [14, 15, 16, 17].
@@
We may also need to take into consideration the boomerang effect, which is the phenomenon in social interaction where attempts to persuade or restrict someone’s freedom of opinion result in the adoption of an opposing view. This effect occurs when persuasive communication happens with an unintended position, which often strengthens the original attitude an individual holds that it sought to alter. Hovland, Janis and Kelly first identified and named the boomerang effect in 1953 [26]. The boomerang effect has been observed in various contexts [27, 28], including consumer behavior, psychological interventions, health policies, and, relevant to our discussion here, political communication. In political contexts, attempts to correct misinformation through mock news articles sometimes lead to increased belief in inaccurate information among certain groups [29].
To study and model the positive persuasion as well as the boomerang effects in social interactions, we introduce a parameter $a\geq 0$ known as the agreement threshold. It is used to determine whether a social interaction results in opinion convergence or divergence. In each time step, we randomly pick an agent from the scale-free network plus one of its neighbors. We count the number of features in which the hyperbolic distance between the opinions of the two agents is smaller than the agreement threshold $a$. If this number is greater than or equal to, say, half of the total number of features, the two agents are said to be “similar”. Otherwise, they are “dissimilar”. We hypothesize that upon interaction, opinions of similar agents are likely to converge whereas those of dissimilar ones are likely to diverge. Besides, an agent’s persuasion power is proportional to the number of neighbors it has. This models the effect that a key opinion leader can change the opinions of his/her followers more effectively than an ordinary individual. Specifically, we state the details of our model below.
@@
    [Opinion initialization] Each agent is characterized by the same finite set of $F$ features =$\mathcal{F}=\{f_{1},\cdots,f_{F}\}$. The opinion of agent $X$ on feature $f\in\mathcal{F}$, denoted by $X_{f}$ is randomly and independently drawn from $\mathcal{N}_{h}(\bar{x},\sigma)=\mathcal{N}_{h}(0,0.5)$, namely, the normal distribution with mean 00 and standard deviation $0.5$ in the hyperbolic distance metric. More precisely, if $r$ is the value of the random variable drawn from $\mathcal{N}_{h}(0,0.5)$, then the opinion is set to $v$ with $d_{\text{h}}(0,v)=r$. Besides, $v\geq 0$ if $r\geq 0$; and $v<0$ if $r<0$.
@@
      [The case of similar agents] If the two agents are similar, namely, the hyperbolic distances between at least $\kappa=1/2$ of their opinions are less than the agreement threshold $a$, then $X_{f}$, the opinion of feature $f$ for agent $X$, changes to $\tilde{X}_{f}$ where
      $d_{\text{h}}(\tilde{X}_{f},X_{f})=\frac{\mu c_{Y}d_{\text{h}}(X_{f},Y_{f})}{c_{X}+c_{Y}}$ (2a)
      with $\operatorname{sgn}(\tilde{X}_{f}-X_{f})=\operatorname{sgn}(Y_{f}-X_{f})$. Here, $c_{X}$ and $c_{Y}$ are the degrees of nodes $X$ and $Y$, respectively. In addition, $\mu>0$ is a parameter modeling the effectiveness of persuasion. Likewise, $Y_{f}$ changes to $\tilde{Y}_{f}$ where
      $d_{\text{h}}(\tilde{Y}_{f},Y_{f})=\frac{\mu c_{X}d_{\text{h}}(X_{f},Y_{f})}{c_{X}+c_{Y}}$ (2b)
      and $\operatorname{sgn}(\tilde{Y}_{f}-Y_{f})=\operatorname{sgn}(X_{f}-Y_{f})$. In other words, the hyperbolic distance of the new opinions of the selected feature trait between the two agents decreases upon interaction.
@@
      [The case of dissimilar agents] If the two agents are dissimilar (that is, they are not similar), then we randomly pick a feature $f$ among those with $d_{\text{h}}(X_{f},Y_{f})\geq a$. The value of $X_{f}$ changes to the unique $\tilde{X}_{f}$ satisfying Eq. (2a) and $\operatorname{sgn}(\tilde{X}_{f}-X_{f})=\operatorname{sgn}(X_{f}-Y_{f})$. Similarly, $Y_{f}$ changes to the unique $\tilde{Y}_{f}$ satisfying Eq. (2b) and $\operatorname{sgn}(\tilde{Y}_{f}-Y_{f})=\operatorname{sgn}(Y_{f}-X_{f})$. Thus, the hyperbolic distance of the new opinions of the selected feature trait between the two agents increases upon interaction.
@@
Note that we use the probability distribution $\mathcal{N}_{h}(0,0.5)$ to initialize opinions to avoid opinion bias as well as to give an initial opinion diversity in our initial ensemble. Last but not least, we remark that our evolution rules do not change the nature of the agent pairs picked in Step 3 in the sense that they are similar (dissimilar) before and after the interaction. Nevertheless, the nature of each of the picked agents with its other neighbors may change after the interaction.

### v2

# Modified Axelrod Model Showing Opinion Convergence And Polarization In Scale-Free Networks

@@ Abstract
Axelrod model is an opinion dynamics model such that each agent on a square lattice has a finite number of possible nominal opinions on a finite number of issues that are usually called features in the field. Moreover, its dynamics between two agents is assimilative in the sense that the number of agreeing features between them never decreases upon interaction. Here we modify this model to study opinion convergence, polarization and more importantly to find ways to reduce opinion polarization in an already polarized population. We do so by changing or adding several elements from complex network and continuous opinion dynamics research. First, we put agents in a scale-free network. Second, we adopt the bounded confidence model by representing our agent’s opinions by numbers in [1,1][-1,1] those distances follow the standard Euclidean metric. Third, our rules allow both convergence and divergence of their resultant opinions after a pair of agents interacts. As a result, our modified model offers a more comprehensive exploration of opinion dynamics. Computer simulation results of our model show scaling behavior and a notable trend in opinion polarization on all features in the majority of reasonable simulation parameters. To mitigate this polarization, we introduce empathetic agents that work actively to reduce opinion differences. However, our findings indicate limited success in the approach for the most effective way is to change the behavior of a significant portion of highly connected agents. This research contributes to the understanding of opinion dynamics within society and highlights the nuanced complexities that arise when considering factors such as network structure and continuous opinion values. Our results prompt further exploration and open avenues for future investigations into effective methods of reducing opinion polarization.
Axelrod model, bounded confidence model, opinion dynamics, opinion polarization, scale-free network
@@ I Introduction
In socio-physics, opinion dynamics of individuals or groups is modeled via interactions among them. By investigating the emergence and convergence of individual opinions into collective opinions, the socio-physical approach aims to shed light on the underlying mechanisms. In particular, sociophysicists wanted to understand the cause as well as the evolution of opinion formation and dissemination through social influence and social networks. They also investigate and analyze the conditions that lead to consensus or polarization, and the factors governing the resulting distributions of opinions [1].
@@
Numerous models of opinion dynamics have been proposed throughout the years. They include the Sznajd model [2, 3], Hegselmann-Krause (H-K) model [4], Galam models [5], Axelrod model [6], Deffuant-Weisbuch model [7, 8] and their variants [9, 10, 11, 12, 13, 9]. Our research here focuses on the Axelrod model. In this model, individuals that we call agents from now on are located on a square lattice. The opinion of each agent, which takes on a discrete set of nominal values, evolves by interacting with neighboring agents according to the set of rules to be described in Sec. II below. In particular, Axelrod found that after a sufficiently long time, all agents on the lattice tend to share the same traits for all features provided that the number of features is sufficiently large or the set of all possible opinions is sufficiently small. Otherwise, multiple regions of different homogeneous opinions on most features is observed in the long run [6]. Castellano et al. found that by varying the size of the discrete possible opinion set, Axelrod model shows the classical order-disorder phase transition on culture fragmentation when each agent has more than two features [14]. Besides, Klemm et al. showed that in the thermodynamic limit, adding an arbitrarily small noise rate to the Axelrod model leads to opinion fragmentation [15]. Opinions in this model are nominal numbers and its evolution rules on a pair of agents are assimilative in the sense that the number of differing opinions between them does not increase immediately after they interact. Thus, this model fails to capture the increasing divided and polarized political as well as ideological stances in numerous countries and societies in our contemporary world though modeling realistic opinions was not the original intention of the Axelrod model.
Various groups of social scientists have proposed reasons for this fragmentation and polarization in the real world [16, 17, 18, 19, 20]. For example, Ashraf and Galor argued that genetic diversity, shaped by the migration of humans out of Africa, is a fundamental determinant of cultural heterogeneity [16]. Dandekar et al. tried to explain the polarization by opinion formation [17].
@@
Socio-physical models could offer a complementary approach to understanding the dynamics of opinion fragmentation. Different scholars have tried different simulation models attempting to capture the fragmentation behaviors in society. See, for example, the review by Sîrbu et al. [9]. Could the Axelrod model be modified to give fragmented and polarized opinions or to incorporate cultural diversity in the long time limit? This is the first question we are going to address. Surely, various authors have studied this issue over the years with partial success. For example, MacCarron et al. considered a more complicated interaction rule and analyzed the clustering behaviors. They showed that by decreasing a certain tunable parameter of the model known as the agreement threshold, agents on the lattice will no longer be homogeneous in the sense that they have more than one set of features in the long run [21]. Radilo-Díaz et al. found that for an initial state with small cultural variability, introducing a small amount of rejection dynamics to the Axelrod model is already sufficient to drive the system eventually to a heterogeneous state [22]. Along another line, Battiston et al. proposed that the interactions in real society are carried out in multiple layers, which correspond to different interests or topics. By considering this multiplicity in their model, they qualitatively presented new dynamical behavior that leads to the emergence and stability of cultural diversity final state [23]. Törnberg proposed a variation of the Axelrod model in which agents have an additional static opinion denoting, say, their partisan or cultural affiliation. Moreover, they can interact nonlocally. As expected, dynamic opinions of agents with the same additional static opinion converge. Interestingly, dynamic opinions of agents with different additional static opinion can be very different. This so-called sorting behavior appears in spite of the absence of repulsive dynamics [24]. Worth noticing that MacCarron et al. [21], Battiston et al. [23] as well as Törnberg [24] considered discrete feature traits. In contrast, Amorim [25] modified the Axelrod model by aligning it with Bourdieu’s theory on the relation between an individual’s taste and the social environment.
This paper introduces a socio-physical model of opinion dynamics that builds upon the Axelrod model, incorporating continuous attitude change among agents within a scale-free network using idea borrowed from the bounded confidence model. First, agents are placed in a more realistic scale-free network of social interaction known as the Barabási–Albert (B-A) network [26]. Second, instead of using a discrete opinion set, we adopt the bounded confidence model [7, 4, 8, 27] by allowing agent’s opinions to take on (continuous) values in the closed interval [1,1][-1,1] under the standard Euclidean distance metric. Third, interaction between a pair of agent may lead to either opinion convergence or divergence. As far as we know, this is the first opinion evolution model that incorporates both assimilative and repulsive dynamics into the Axelrod model with a bounded confidence favor.
Certainly, opinion dynamics studies on scale-free network are not new. For example, Bartolozzi et al. discovered intermittence in a probabilistic two-state model of opinion evolution on the B-A network [28]. Jacobmeier studied evolution of discrete cardinal opinion evolution for agents in the B-A network using assimilative dynamics [29]. More importantly, Dinkelberg et al. found that the long term behavior of a modified Axelrod model with discrete cardinal opinions is insensitive to network topology [30]. Similarly, various groups have studied opinion dynamics using the bounded confidence model [7, 4, 8, 27]. (See, for example, the reviews in Refs. [31] and [32] for more details.) In particular, Deffuant et al. put forward several agent-based models with continuous opinion change such that the involved agents have limited ranges of acceptance and non-commitment, where only the impact of assimilation is taken into account [7, 8]. Fortunato et al. showed that the eventual states of the combination of the H-K model and the bounded confidence model with assimilative interaction only may consist of multiple homogeneous opinion clustering regions [33]. Lorenz investigated the opinion dynamics of agents placed on $d$-dimensional cubes or simplexes using bounded confidence model with assimilative interaction [34]. Along a different line, Huet and Deffuant proposed a bounded confidence model with major and minor feature traits whose interaction can be assimilative or repulsive [35]. Moreover, Chau et al. [13] proposed a model for opinion formation which is an extension of the Jager-Amblard (J-A) model [27] under the framework of bounded confidence. Chau et al.’s model incorporates assimilation and contrast, which influence agents’ opinion acceptance or rejection. More importantly, it shows opinion clustering around extremes and/or moderate. Furthermore, with slow variation of model parameters, Chau et al. observed first-order phase transition in the distribution of opinion [13]. Recently, Liu et al. researched on the complicated dynamics of the bounded confidence opinion evolution with assimilative and repulsive interactions for agents arranged on a ring [36]. Last but not least, we remark that the works of Jacobmeier [29], Lorenz [34] and Dinkelberg et al. [30] mentioned earlier mixed the Axelrod model with the bounded confidence model. But unlike ours, they did not include repulsive dynamics.
Through computer simulation, we show that agents in our model gradually develop extreme opinions. These findings are consistent with observed social behavior and hence give novel insights into the factors influencing collective human opinions. Scaling is also observed in the opinion cluster size distribution whose precise definition will be given later in the text.
Most importantly, we also attempted to address our second question, namely, the polarization problem that is so urgently needed in our currently highly divided and politicized society by modifying our model in this paper. The intuition is that higher openness to ideas from the opposite camp may lead to mutually understanding and increase their willingness to seek common grounds. Therefore, we set a small portion of agents that are empathetic by actively trying to reduce opinion differences. (Note that introducing heterogeneous agents has been done before [8], but setting empathetic agents in this way is new.) Nonetheless, this approach can only economically reduce opinion polarization of an already polarized population by making a significant portion of those highly connected agents empathetic. This work is an extension of a short project done by XZ [37] under the supervision of HFC, the teacher of the course “Data Analysis And Modeling In Physics”.
@@ II The Axelrod Model
One of the most frequently used ways of measuring individual opinions in a society is to collect data from attitudinal surveys over a long period of time. It is instructive to model the attitude evolution of individuals and interest groups. One such model was proposed by Axelrod [6]. In the Axelrod model, agents are placed on a $N\times N$ square lattice with closed boundary conditions. An agent is conveniently labeled by its position $(i,j)$ on the lattice. Each agent is characterized by the same fixed set of $F$ features $\mathcal{F}=\{f_{1},f_{2},\cdots,f_{F}\}$. Each feature takes only values from a fixed set of traits. For simplicity, each $f_{i}\in\mathcal{F}$ is taken from the same set of $q$ traits labeled $\{0,1,\dots,q-1\}$. As a result, each agent has $F^{q}$ possible internal states. For example, one of the features could be the primary language spoken by the agent, and the corresponding set of traits is the set of all contemporary languages. Another possible feature is one’s religious belief. Each feature of an agent is randomly assigned in the beginning. Furthermore, the system evolves for several time steps as follows.
@@
Under the Axelrod model, agents’ opinions converge provided that $F$ is sufficiently small or $q$ is sufficiently large.
@@ III Deficiencies Of The Axelrod Model And Our Remedy
The Axelrod model is simple and useful in opinion dynamics. However, it has several deficiencies even though being realistic was not the original goal of the Axelrod model. We now point them out and report ways to modify the model so as to produce an increasingly polarized opinion distribution that is robust under small changes in tunable parameters.
@@ III.1 Network Topology
In the Axelrod model, agents form a two-dimensional square lattice. This network topology is artificial. More importantly, it is not compatible with the fact that many social structures can be described by scale-free networks [38]. In simple terms, all but those agents in the lattice boundary in the Axelrod model have equal strength in influencing others or being influenced by others. In contrast, scale-free network topology can be used to model both social and parasocial interactions in the real world, including interactions between social media influencers and their followers. Thus, we propose putting agents in an (undirected) scale-free network in which only neighboring agents are allowed to interact directly. For simplicity, we generate such a network using the B-A model [26] as it is one of the most studied and used scale-free networks. Besides, it has the additional advantage that it can be efficiently generated. In this study, we simplify matters by considering non-evolving networks. It is instructive to consider evolving networks in future studies.
@@ III.2 Feature Traits
In the Axelrod model, every agent is described by a set of features, and each feature can have several possible discrete nominal traits [6]. While there are features that are naturally described by discrete nominal traits, the Axelrod model is incomplete as traits of a feature need not be discrete or nominal. For example, one of the features could be the agent’s fluency in a particular spoken language. Here, the possible traits are more naturally described by a continuous real number than discrete nominal states. More importantly, using continuous traits is a natural platform to model the possibly increasingly polarized opinion distribution. We, therefore, modify the model by replacing the discrete set of traits with a real number representing the strength of the agent’s feature. We also call this type of continuous feature traits opinions of the agent.
How could we take care of the increasingly polarized opinion distribution? A simple and effective way to address this issue is to borrow the idea of the bounded confidence model [7, 4, 8, 27] by representing opinions using a real number in a bounded interval under the Euclidean metric. For simplicity, we set the possible opinions to be a real number in the interval [1,1][-1,1]. Besides, we denote the distance between two opinions values $p$ and $q$ by
d(p,q)=|pq|.d(p,q)=|p-q|. (1)
As for the initial values of each trait, instead of drawing from a uniform distribution in [1,1][-1,1], we use a more realistic truncated normal distribution with respect to the Euclidean distance metric. Its detailed construction can be found in upcoming subsection.
@@ III.3 Interaction Rules Among Agents
Recall that agents change their traits according to rule 4 in the Axelrod model. This rule is assimilative and hence can only increase the cultural similarity between the chosen pair of agents. In other words, rule 4 is too cohesive. Besides, it is not consistent with the observation that cultural diversity among groups of people appears even in apparently culturally homogeneous societies [16, 17, 18, 19].
@@
We may also need to take repulsive dynamics into consideration. The existence of repulsive dynamics and its psychological origin (if it exits) are current research topics. One possibility is that the effect is real. Supporters of this camp believe that social interaction where attempts to persuade or restrict someone’s freedom of opinion could result in the adoption of an opposing view. This effect occurs when persuasive communication happens in an unintended position, which often strengthens the original attitude an individual holds that it sought to alter. Hovland et al. first identified and named it the boomerang effect [39]. The boomerang effect has been observed in various contexts [40, 41], including consumer behavior, psychological interventions, health policies, and, relevant to our discussion here, political communication. In political contexts, attempts to correct misinformation through mock news articles sometimes lead to increased belief in inaccurate information among certain groups [42]. Nevertheless, we need to the caution here as the evidence on the existence of the boomerang effect is controversial and mixed. Furthermore, this effect may not be systematic. For example, Schultz et al. observed it in a field experiment in promoting household energy conservation through the use of different normative messages. Moreover, boomerang effect was eliminated by adding an injunctive message [43]. Similar conclusions were drawn in the follow up study by Verkooijen et al. [44]. In addition, Liu et al. found that overinvestment in corporate social responsibility had a boomerang effect on the shareholder value of a company after announcing a recall. The effect is more evident when institutional ownership is low or when customer awareness is high [45]. On the other hand, Takács et al. failed to observe it in their recent study using the University of Groningen students as participants [46]. Bail et al. discovered that exposing Democrats (Republicans) with conservative (liberal) social media messages make their views more liberal (conservative). However, the change is statistically insignificant [47]. Another possibility is that boomerang effect is neither real nor systematic. At least, agents tend to be less affected by ideas from the opposite camp [24]. If it is so, what is the underlying mechanism for opinion polarization? Unfortunately, no definitive answer has been identified.
Here we move on by adopting an empirical approach. That is to say, we model the repulsive dynamics as if it were coming from the boomerang effect without considering its psychological origin. Our goal is to reproduce the polarization seen in the real world and study if there is anything we can do to reduce it. Specifically, to study and model the positive persuasion as well as the repulsion effects in social interactions, we introduce a parameter $a\geq 0$ known as the agreement threshold. It is used to determine whether a social interaction results in opinion convergence or divergence. In each time step, we randomly pick an agent from the scale-free network plus one of its neighbors. We count the number of features in which the Euclidean distance between the opinions of the two agents is smaller than the agreement threshold $a$. If this number is greater than or equal to, say, half of the total number of features, the two agents are said to be “similar”. Otherwise, they are “dissimilar”. We hypothesize that upon interaction, opinions of similar agents are likely to converge whereas those of dissimilar ones are likely to diverge. Besides, an agent’s persuasion power is proportional to the number of neighbors it has. This models the asymmetric influence between agents of different connectivities. This hypothesis is consistent with a number of findings in the literature on the persuasion power of social media influencers. They include the meta-analysis by Pan et al. in which they found that the characteristics of a social media influencer, including the number of followers, is positively correlated with the purchasing behavior of his/her followers [48]. The report by Conde and Casais showed the positive correlation of the influencing power with the perceived popularity and opinion leadership from a sample of 140 Portuguese social media influencers [49]. The work by Peter and Muth who discovered that social media influencers were increasingly affecting German youth on political issues. The effects ranged from amplifying the effects of existing opinions to opinion formation as well as to change voting intentions [50]. Last but not least, the work by Liu and Zheng who argued that the influence between social media influencers and their followers is mutual when evaluating brand credibility and purchase intention. They further found that the influence could be repulsive sometimes [51]. Nevertheless, we have to be caution that little is known on how the followers directly affecting the positions of social media influencers. It is instructive and timely to investigate the validity of the hypothesis that an agent’s persuasion power is proportional to its number of neighbors in the field experiments.
After the above preparation and discussion, we now state the details of our model below.
@@
    [Opinion initialization] Each agent is characterized by the same finite set of $F$ features $\mathcal{F}=\{f_{1},\cdots,f_{F}\}$. The opinion of agent $X$ on feature $f\in\mathcal{F}$, denoted by $X_{f}$ is randomly and independently drawn from $\mathcal{N}_{t}(\bar{x},\sigma)=\mathcal{N}_{t}(0,0.25)$, namely, the truncated normal distribution with mean 0 and standard deviation $0.25$ in the Euclidean distance metric. More precisely, if $r$ is the value of the random variable drawn from $\mathcal{N}_{t}(0,0.25)$, then the opinion is set to $r$ if $|r|\leq 1$. And the opinion is set to 11 (1)(-1) if $r>1$ (r<1r<-1).
@@
      [The case of similar agents] If the two agents are similar, namely, the Euclidean distances between at least $\kappa=1/2$ of their opinions are less than the agreement threshold $a$, then their opinions of feature $f$ move closer to each other in the sense that the Euclidean distance of the new opinions of the selected feature trait between the two agents decreases upon interaction. Specifically, $X_{f}$, the opinion of feature $f$ for agent $X$, changes to $\tilde{X}_{f}$ where {subequations}
      $\tilde{X}_{f}=X_{f}-\frac{\mu c_{Y}|X_{f}-Y_{f}|}{c_{X}+c_{Y}}.$ (2)
      Here, $c_{X}$ and $c_{Y}$ are the degrees of nodes $X$ and $Y$, respectively. In addition, $\mu>0$ is a parameter modeling the effectiveness of persuasion. Note that we borrow the preferential attachment idea in the B-A network [26] in this rule so as to model the effective persuasion power of social media influencers. Likewise, $Y_{f}$ changes to $\tilde{Y}_{f}$ where
      $\tilde{Y}_{f}=Y_{f}-\frac{\mu c_{X}|Y_{f}-X_{f}|}{c_{X}+c_{Y}}$ (3)
@@
      [The case of dissimilar agents] If the two agents are dissimilar (that is, they are not similar), then we randomly pick a feature $f$ among those with $d(X_{f},Y_{f})\geq a$. The values of $X_{f}$ and $Y_{f}$ change to $\tilde{X}_{f}$ and $\tilde{Y}_{f}$ according to {subequations}
[…]
      and
[…]
      respectively. Here
[…]
      Clearly, the Euclidean distance of the new opinions of the selected feature trait between the two agents increases or remains unchanged upon interaction.
@@
Note that we use the probability distribution $\mathcal{N}_{t}(0,0.5)$ to initialize opinions to avoid opinion bias as well as to give an initial opinion diversity in our initial ensemble. In addition, our evolution rules do not change the nature of the agent pairs picked in Step 3 in the sense that they are similar (dissimilar) before and after the interaction. Nevertheless, the nature of each of the picked agents with its other neighbors may change after the interaction. Last but not least, we stress that for all the subsequent simulations that study the eventual system behavior, the average number of interactions an agent has with its neighbors $t$ is chosen to be large enough for the system to approach equilibration.